let (7, -5) be a point on the terminal side of feta. find the exact values of cos feta, sec feta, and cot feta.
Answer by lwsshak3(6489)

let (7, -5) be a point on the terminal side of feta. find the exact values of cos feta, sec feta, and cot feta.
**
use x for feta
Given point places reference angle in quadrant IV
You are working with a right triangle with:
opp side=-5
adj side=7
hypotenuse=√(5^2+7^2)=√(25+49)=√74
..
cosx=7/√74
secx=√74/7
cotx=-7/5